Calculate the cell potential of the galvanic cell at 298К consists of two silver electrodes immersed into two solutions having с1(Ag+)= 10 –2mol/L and с2(Ag+)= 10 –1mol/L.


1
Expert's answer
2021-11-13T01:33:47-0500

Cell potential E0 = -0.059×log("\frac{c_1}{c_2}" )

E0 = -0.059× log("\frac{0.01}{0.1}" )

E0 = 0.059V